1) What is the future value of $8,000 invested today and held for 15 years at 8.5 percent compounded annually?

2) Your grandparents just gave you a gift of $3,000. You are investing this money for 10 years at 3 percent simple interest. How much money will you have at the end of the 6 years?

3) Today, you deposit $2,500 in a bank account that pays 3.6 percent simple interest. How much interest will you earn over the next 5 years?
Answer #1

1). FV = PV * (1 + r)n

= $8,000 * (1 + 0.085)15 = $8,000 * 3.3997 = $27,197.94

2). Simple Interest Earned = P * R * T = $3,000 * 0.03 * 6 = $540

Value at the end of 6 years = Value of investment + Interest earned = $3,000 + $540 = $3,540

3). Simple Interest Earned = P * R * T = $2,500 * 0.036 * 5 = $450